OpenStudy (anonymous):

How can I show the series from 0 to infinity (1/n+3)-(1/n+1) converges?

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty}\frac{ 1 }{ n+3}-\frac{ 1 }{ n+1}\]

OpenStudy (anonymous):

Have you tried the comparison test

OpenStudy (anonymous):

I have. But I'm confused because the series is always negative. I'm not sure what I could compare it to.

OpenStudy (kainui):

You should consider writing out a few terms of the series and see if you notice any patterns.

OpenStudy (anonymous):

I'll let kai handle this one xD

OpenStudy (kainui):

Nope, that's all I'm saying lol I'm busy atm

OpenStudy (anonymous):

hmmmm so its getting increasingly less negative? Looks like it converges to 0 maybe?

OpenStudy (anonymous):

But I'm still not sure what I could compare it to...

OpenStudy (anonymous):

It's a telescoping series, just do what kai said and you should get it ;d

OpenStudy (anonymous):

Maybe 1/n^2

OpenStudy (anonymous):

I thought about that but 1/n^2 is always larger than this function.

OpenStudy (kainui):

You don't compare it to anything since you can exactly evaluate it.

OpenStudy (anonymous):

Oh wait, I said that wrong.

OpenStudy (anonymous):

Exactly evaluate it? how?

myininaya (myininaya):

try writing out the first few terms as @Kainui suggested

OpenStudy (anonymous):

^

myininaya (myininaya):

you should see a lot of zeroing out

OpenStudy (anonymous):

For the first three terms I got -.66, -.25, and -.133

myininaya (myininaya):

i would leave them infractional form

myininaya (myininaya):

in fractional *

OpenStudy (anonymous):

I see its getting closer to zero....But thats not enough right?

OpenStudy (anonymous):

Oh ok. I'll do that.

OpenStudy (anonymous):

-2/3, -1/4,-2/15...

OpenStudy (anonymous):

I'm sorry, im still not sure how that helps me =/

myininaya (myininaya):

well that doesn't help

myininaya (myininaya):

it looks like you combined the fractions

OpenStudy (kainui):

\[\LARGE \sum_{n=0}^\infty \frac{1}{n+3}-\sum_{n=0}^\infty \frac{1}{n+1}\] We can see that by replacing n with k+2 that it looks similar for the second summation \[\LARGE \sum_{n=0}^\infty \frac{1}{n+3}-\sum_{k+2=0}^\infty \frac{1}{k+2+1}\]\[\LARGE \sum_{n=0}^\infty \frac{1}{n+3}-\sum_{k=-2}^\infty \frac{1}{k+3}\] Now we pull out the -2 and -1 terms leaving us at k=0 so that it matches the other one: \[\LARGE \sum_{n=0}^\infty \frac{1}{n+3}-\frac{1}{1}-\frac{1}{2}-\sum_{k=0}^\infty \frac{1}{k+3}\] but since k and n are just dummy variables, these both represent the exact same summation. So they cancel out leaving us with just a couple terms. Fancy yeah?

OpenStudy (anonymous):

ohhhh I'm getting that!!

myininaya (myininaya):

Here an example: \[\sum_{n=1}^{\infty}(\frac{1}{n}-\frac{1}{n+1}) \\ (\frac{1}{1}-\frac{1}{1+1})+(\frac{1}{2}-\frac{1}{1+2})+(\frac{1}{3}-\frac{1}{1+3})+(\frac{1}{4}-\frac{1}{4+1}) +\cdots \\ (\frac{1}{1}-\cancel{\frac{1}{1+1}})+(\cancel{\frac{1}{2}}-\cancel{\frac{1}{1+2}})+(\cancel{\frac{1}{3}}-\cancel{\frac{1}{1+3}})+(\cancel{\frac{1}{4}}-\frac{1}{4+1}) +\cdots \\ =\frac{1}{1}=1 \] I didn't put a little cancelly thing over the 1/5 part but it does cancel

OpenStudy (anonymous):

Ok I think ive got it. Thanks to all that helped out!!

myininaya (myininaya):

I like the what @Kainui wrote it

myininaya (myininaya):

way*

OpenStudy (anonymous):

Yeah me to... Never would have thought to do it that way.

OpenStudy (kainui):

Yeah, it's a lot like "u-substitution" when solving an integral. Once you know about this way, if you can transform the sum into the other sum by a simple little substitution like this you can sometimes get them to work out pretty nicely like this. =D